OpenStudy (anonymous):

a jogger ran at a rate of 5.4 miles per hour find the joggers exact rate feet per minute 1 mile-5280 feet

OpenStudy (anonymous):

convert 5.4 miles to feet=28512

OpenStudy (anx):

1 mine = 5280 feet 1 hour = 60 min so 5.4 (5280) = 28512 feet 28512 / 60 = 475.2 feet per hour 475.2 /60 = 7.92 or 8 feet per minute

OpenStudy (anonymous):

and hours to minutes, 1 hr=60 min

OpenStudy (anonymous):

divide 28512 by 60, =475.2

OpenStudy (anonymous):

@anx that is feet per second

OpenStudy (anx):

oops, went a step to far...its 28512 / 60 = 475.2 feet per minute.

OpenStudy (anonymous):

i have to show how exactly you got 5.4 miles to feet

OpenStudy (anonymous):

you took it too far

OpenStudy (anx):

5.4 miles (5280 feet) there are 5280 feet in a mile you have 5.4 miles 5.4*5280 will give you how many feet you have.

OpenStudy (anonymous):



OpenStudy (anonymous):

thankyou both!